# DisCoRD: Discrete Tokens to Continuous Motion via Rectified Flow Decoding

### v1

@@ Abstract
Human motion, inherently continuous and dynamic, presents significant challenges for generative models. Despite their dominance, discrete quantization methods, such as VQ-VAEs, suffer from inherent limitations, including restricted expressiveness and frame-wise noise artifacts. Continuous approaches, while producing smoother and more natural motions, often falter due to high-dimensional complexity and limited training data. To resolve this “discord” between discrete and continuous representations, we introduce DisCoRD: Discrete Tokens to Continuous Motion via Rectified Flow Decoding, a novel method that decodes discrete motion tokens into continuous motion through rectified flow. By employing an iterative refinement process in the continuous space, DisCoRD captures fine-grained dynamics and ensures smoother and more natural motions. Compatible with any discrete-based framework, our method enhances naturalness without compromising faithfulness to the conditioning signals. Extensive evaluations demonstrate that DisCoRD achieves state-of-the-art performance, with FID of 0.032 on HumanML3D and 0.169 on KIT-ML. These results solidify DisCoRD as a robust solution for bridging the divide between discrete efficiency and continuous realism. Our project page is available at: Project Page.
@@ 1 Introduction
Human motion generation controlled by diverse signals has become a pivotal area in computer vision, driven by its vast applications in virtual reality to animation, gaming, and human-computer interaction. The ability to generate realistic human motions that are precisely aligned with input conditions—such as textual descriptions [44, 13, 55, 11], human speech [58, 25, 27], or even music [20, 40, 9]—is essential for creating immersive and interactive experiences. Two critical qualities define the success of such systems [49]: faithfulness, ensuring that the generated motion accurately reflects the conditioning signal, and naturalness, producing smooth and lifelike motions that are comfortable and convincing to human observers. A deficit in faithfulness can produce motions that are misaligned or irrelevant to the intended context, while even slight unnaturalness can disrupt user immersion, creating an uncanny valley effect [32].
@@
Since human motion is inherently continuous, human motion generation based on continuous representations is inherently well-suited for producing smooth and natural motion [61, 44, 55, 33]. However, due to the high dimensionality of continuous representation, they often encounter challenges with cross-modal mapping ambiguity [56, 42] which can result in low faithfulness. This issue becomes especially pronounced in data-constrained settings, such as motion capture datasets [30], where limited examples lead to difficulty of learning consistent mappings between signals and motions. On the other hand, discrete quantization methods [13, 36, 60] utilize motion VQ-VAEs [48] to discretize motion representation, simplifying the learning of high-dimensional data mappings by reformulating it as a classification task. This discretization enables more efficient learning and can be particularly beneficial when dealing with limited data, improving faithfulness. However, motion VQ-VAEs face two main challenges: under-reconstruction, caused by representing continuous motion with a finite set of tokens, and frame-wise noise, caused when decoding these discrete representations. Under-reconstruction can lead to loss of fine-grained motion details, essential for dynamic motion, and frame-wise noise can introduce unnatural artifacts, degrading the naturalness of generated motions and diminishing user immersion.
@@
Figure 1: Concept of DisCoRD. Discrete quantization methods encode multiple motions into a single quantized representation. While existing methods deterministically decode from this quantized representation, DisCoRD iteratively decodes the discrete latent in a continuous space to recover the inherent continuity and dynamism of motion. To assess the gap between reconstructed and real motion, prior work primarily used FID as the metric. Here, we additionally propose symmetric Jerk Percentage Error (sJPE) to evaluate the differences in naturalness between reconstructed and real motion.
@@
In this paper, we propose DisCoRD, a novel approach that combines the strengths of continuous and discrete methods to achieve natural and faithful human motion generation. Our key insight is to leverage the strong faithfulness of discrete generation methods [36, 13] by employing score-based models [14, 41, 26, 24] to translate pretrained discrete tokens back into raw motion space, enhancing naturalness while preserving precise alignment with conditioning signals.
Our approach offers two primary advantages over traditional discrete decoding methods, shown in Figure 1. First, by replacing the deterministic, feed-forward decoding of standard quantization techniques with stochastic, iterative refinement with score-based models, we can restore lost details from the quantization process [5, 64] which is not possible with a deterministic decoder. This enables the reconstruction of dynamic and complex movements that were challenging for previous methods. Second, instead of directly decoding discrete tokens, we employ them exclusively as a condition to steer motion generation in the raw-motion space, effectively reducing fine-grained noise and producing smoother, more natural motion outputs.
@@
While our method generates natural motions, evaluating the naturalness of reconstructed motions remains challenging as traditional metrics like MPJPE do not align well with human perception [8], and FID fails to capture subtle, frame-wise noise, shown in Figure 3. This limitation reduces the effectiveness of these metrics in assessing naturalness, particularly in cases of under-reconstruction and frame-wise noise. To address this, we introduce a novel metric: the symmetric Jerk Percentage Error (sJPE), designed to evaluate reconstructed motion by ensuring it is free from under-reconstruction while also minimizing fine-grained noise. Our experiments demonstrate the effectiveness of DisCoRD in enhancing sample-wise naturalness without suffering from under-reconstruction.
Additionally, our method is adaptable to any discrete-based motion generation framework, independent of conditioning signals, to improve performance. Extensive experiments across diverse tasks—text-to-motion generation, co-speech gesture generation, and music-to-dance generation—demonstrate that our approach integrates seamlessly with various existing methods, achieving state-of-the-art naturalness in most cases according to both standard evaluation metrics and our proposed metric, sJPE.
Overall, our contributions are as follows:
@@
    We introduce DisCoRD, a novel method for decoding discrete tokens in continuous motion space, enhancing the naturalness of generated motions, while maintaining faithfulness.
@@
    We propose a novel evaluation scheme, the symmetric Jerk Percentage Error, designed to evaluate both under-reconstruction and frame-wise noise, which are often overlooked but critical for motion generation.
@@ 2 Related Work
The goal of human motion generation is to generate natural motions that adhere to control signals. Recent advancements in this field can be broadly categorized into two main approaches: motion generation from continuous representations, where models directly regress on continuous values, and motion generation from discrete representations, where motion is quantized into discrete tokens, transforming the generation task into a classification problem in the discrete domain.
Motion Generation from Continuous Representation.
Motion generation using continuous representations can be achieved through various approaches. Early methods for signal-to-motion generation primarily focused on mapping control signals to motions within a continuous representation space [10, 2, 34, 43, 22], often leveraging the Variational Autoencoder (VAE) architecture [17] or utilizing CLIP features [38]. More recently, with the success of score-based models [14, 41], new advancements have emerged in generating motion through continuous representations [45, 55, 62, 6, 61, 59, 46, 23, 21, 29]. These approaches aim to model the inherently continuous nature of motion using a continuous representation space, making them well-suited to the task of motion generation. However, due to the data-constrained nature of the current motion generation tasks, the complexity of the continous representation space often makes it challenging to establish a reliable cross-modal mappings between control signals and motions, resulting in suboptimal performance.
Motion Generation from Discrete Representation.
Recently, to simplify the complex mapping between control signals and motions, some methods have reformulated the generation task as a discrete classification problem, achieving notable performance in motion generation [36, 60, 12, 13, 35, 16, 40, 57, 27, 58, 25]. These approaches often employ VQ-VAE [47] to create motion tokens, which are then used to generate motion sequences via token prediction models, such as autoregressive transformers. More recently, discrete diffusion models have been introduced to directly denoise these discrete tokens [7, 28]. While these methods effectively bypass complex signal-to-motion mapping challenges, their inherent characteristics—such as information loss and discreteness—frequently result in unnatural artifacts, including under-reconstruction and frame-wise noise.
Our method combines the inherent naturalness of continuous approaches with the signal faithfulness of discrete methods, achieving both natural and well-conditioned motion generation within the data-constrained motion generation task.
@@ 3 Method
In this section, we propose DisCoRD, a novel method for decoding pretrained discrete representations in the continuous domain using score based models. This allows the generated motions to: 1) capture dynamic, fast-paced motion due to iterative and stochastic nature of score-based models and 2) maintain the natural smoothness of motion by decoding in the raw motion domain, conditioned on discrete tokens. We begin by introducing score-based models, including rectified flow, then present the concept of motion tokenization. This is followed by an explanation of our conditional rectified flow model, conditioning projection module, and training details.
@@ 3.1 Preliminaries
Score-based Models.
Given a data distribution $p_{x}$, score-based models [14, 41] define a forward diffusion process that progressively adds noise to the data $x_{0}\sim p_{x}$, eventually transforming it into $x_{T}\sim\mathcal{N}(0,I)$. Typically, this process can be expressed as:
$x_{t}=\alpha_{t}\cdot x_{0}+\sigma_{t}\cdot x_{T},\ \mathrm{where}\ t\in[0,T].$ (1)
Here, the choice of functions $\alpha_{t}$ and $\sigma_{t}$ determines the trajectory that connects the data and noise. Score based model then learns the reverse diffusion process to recover the original data from noise. Recent works have introduced various parameterizations for $\alpha_{t}$ and $\sigma_{t}$ to enhance the reverse diffusion process. Beyond specific parametrization of $\alpha_{t}$ and $\sigma_{t}$, the iterative nature of score-based models significantly strengthens their capacity to capture complex data variations, particularly excelling in generating high-dimensional data.
@@
Rectified flow [26] is a generative model developed to address the transport mapping problem by employing a flow matching algorithm [26, 1, 24]. Flow matching focuses on constructing a transport map, denoted as $T:\mathbb{R}^{d}\rightarrow\mathbb{R}^{d}$, that effectively maps observations from source distribution $x_{0}\sim\pi_{0}$ on $\mathbb{R}^{d}$ to observations from target distribution $x_{1}\sim\pi_{1}$ on $\mathbb{R}^{d}$. This transport map is formalized via the ordinary differential equation (ODE):
@@
In this context, $v$ is typically referred to as a vector field, while $x_{t}$ represents the forward process parameterized over $t\in[0,1]$. Rectified flow specifies a parametrization of Equation  (1) such that the forward process follows a straight path. This trajectory is expressed as $x_{t}=tx_{1}+(1-t)x_{0}$, with $v$ defined as $(x_{1}-x_{0})$. Since the target $x_{1}$ is unknown during the generation phase, a simple least squares regression problem:
@@
is used to train a causal approximation of $v$, denoted $v_{\theta}$. Once the vector field $v_{\theta}$ has been trained, samples from target distribution $\pi_{1}$ can be generated by solving Equation (2) using an ODE solver, where the initial conditions for the ODE solver comes from source distribution $\pi_{0}$. Due to the nearly straight trajectory that directly bridges the source and target distributions, the rectified flow model effectively enables the efficient transport of samples from readily accessible distributions to the target distribution.
@@ 3.2 DisCoRD
Traditional deterministic feed-forward decoders $\mathcal{D}$ suffer from limited expressiveness and propagate token discreteness into decoded motions, resulting in under-reconstructed and noisy outputs (see Figure 4). To address these issues, we propose decoding pretrained tokens in continuous space by replacing $\mathcal{D}$ with an expressive score-based model. Specifically, we propose a rectified flow model that synthesizes human motion from Gaussian noise, conditioned on frame-wise motion features. These features are extracted by transforming discrete tokens through a conditional projection module, resulting in expressive frame-wise motion representations. The overall pipeline of DisCoRD is depicted in Figure 2.
@@
We approach token decoding as a conditional generation task using rectified flow. Specifically, given a motion data distribution $P_{\mathbf{X}}$, we define the transport process from Gaussian noise $\mathbf{X_{0}}\sim\mathcal{N}(0,I)$ to motion $\mathbf{X_{1}}\sim P_{\mathbf{X}}$, guided by $\mathbf{C}$ which are features extracted from discrete tokens, formulated as:
@@
In contrast to traditional score-based models that generate images or motion [44] from sparse conditions like text, our goal is to decode discrete codebooks into raw motion using dense, frame-wise conditions. We achieve this by concatenating frame-wise features extracted from discrete tokens along the channel dimension, similar to image generation methods [64, 15]. This setup allows each motion frames to be conditioned on per-frame codebook features, enabling more expressive decoding. During the inference stage, features $\mathbf{\hat{C}}$, extracted from tokens by a pretrained token generation model, are iteratively decoded into $\mathbf{\hat{X}_{1}}$ by solving a conditional ODE using the Euler method.
Condition Projection.
To enable our decoder to generate expressive motion, we extract continuous, frame-wise conditioning features from downsampled discrete tokens $\mathbf{Z}=[\mathbf{z}_{1},\dots,\mathbf{z}_{T/q}]$. Rather than applying naive upsampling and linear projection—which would lose temporal information—each token $\mathbf{z}_{t}$ is repeated $q$ times to restore the original temporal resolution. The repeated $q$ tokens are then stacked into a tensor $\mathbf{Z}_{\text{stacked}}\in\mathbb{R}^{(T/q)\times(q\times d_{code})}$ and processed by a linear layer. Subsequently, we unstack the processed tensor to yield a sequence of frame-wise conditioning features, denoted as $\mathbf{C}=[\mathbf{c}_{1},\dots,\mathbf{c}_{T}]$, where each frame is assigned a unique set of temporal features. This method enhances the expressiveness of our frame-wise features, effectively guiding the iterative decoding process for improved motion quality.
@@
In typical score-based models, training is performed over the entire data instance with fixed sequence lengths, such as 196 frames in HumanML3D [11], a standard in motion generation [44, 55]. While this approach improves reconstruction quality in stage 1, our results indicate that these improvements do not translate effectively to generation quality in stage 2. To address this limitation and improve generalization to unseen motion sequences during inference, we train our rectified flow model on sliding windows of motion frames rather than max length spans. Additionally, although conventional U-Net diffusion models often incorporate attention mechanisms to enhance performance [39], we found this strategy to be suboptimal in our context, resulting in a performance degradation.

### v3

@@ Abstract
Human motion is inherently continuous and dynamic, posing significant challenges for generative models. While discrete generation methods are widely used, they suffer from limited expressiveness and frame-wise noise artifacts. In contrast, continuous approaches produce smoother, more natural motion but often struggle to adhere to conditioning signals due to high-dimensional complexity and limited training data. To resolve this “discord” between discrete and continuous representations we introduce DisCoRD: Discrete Tokens to Continuous Motion via Rectified Flow Decoding, a novel method that leverages rectified flow to decode discrete motion tokens in the continuous, raw motion space. Our core idea is to frame token decoding as a conditional generation task, ensuring that DisCoRD captures fine-grained dynamics and achieves smoother, more natural motions. Compatible with any discrete-based framework, our method enhances naturalness without compromising faithfulness to the conditioning signals on diverse settings. Extensive evaluations demonstrate that DisCoRD achieves state-of-the-art performance, with FID of 0.032 on HumanML3D and 0.169 on KIT-ML. These results establish DisCoRD as a robust solution for bridging the divide between discrete efficiency and continuous realism. Code and checkpoints will be released.
@@ 1 Introduction
Human motion generation controlled by diverse signals has become an emerging area in computer vision, driven by its vast applications in virtual reality to animation, gaming, and human-computer interaction. The ability to generate realistic human motions that are precisely aligned with input conditions—such as textual descriptions [51, 12, 60, 11], human speech [62, 28, 30], or even music [25, 46, 10]—is essential for creating immersive and interactive experiences. Two critical qualities define the success of such systems [56]: faithfulness, ensuring that the generated motion accurately reflects the conditioning signal, and naturalness, producing smooth and lifelike motions that are comfortable and convincing to human observers. Deficiencies in faithfulness cause misaligned motions, while slight unnaturalness disrupts immersion and triggers the uncanny valley [36] effect.
@@
Since human motion is inherently continuous, generation based on continuous representations is naturally well-suited for producing smooth and realistic motion [65, 51, 60, 38]. However, due to the high dimensionality of continuous representation, they often encounter challenges with cross-modal mapping ambiguity [61, 49] which can result in low faithfulness. This issue becomes especially pronounced in data-constrained settings, such as motion capture datasets [32], where limited examples lead to difficulty of learning consistent mappings between signals and motions. On the other hand, discrete quantization methods [12, 41, 64] utilize motion VQ-VAEs [55] to discretize motion representation, simplifying the learning of high-dimensional data mappings by reformulating it as a classification task. This discretization enables more efficient learning and can be particularly beneficial when dealing with limited data, improving faithfulness [12, 40, 41]. However, motion VQ-VAEs face two main challenges. First, under-reconstruction occurs when fine-grained motion details, which are essential for generating dynamic movements, are lost during token discretization. Second, frame-wise noise arises from directly decoding discrete tokens, introducing unnatural artifacts that disrupt motion smoothness and diminish user immersion. These challenges make it difficult to generate motion that is both smooth and natural while maintaining high faithfulness.
@@
Figure 1: Concept of DisCoRD. Discrete quantization methods encode multiple motions into a single quantized representation. While existing methods directly decode from this quantized representation, DisCoRD iteratively decodes the discrete latent in a continuous space to recover the inherent continuity and dynamism of motion. To assess the gap between reconstructed and real motion, prior work primarily used FID as the metric. Here, we additionally propose symmetric Jerk Percentage Error (sJPE) to evaluate the differences in naturalness between reconstructed and real motion.
@@
In this paper, we propose DisCoRD, a novel approach that bridges discrete and continuous motion generation to achieve both faithfulness and naturalness. Our key insight is to leverage the strong faithfulness of discrete generation methods [64, 12, 40] by utilizing rectified flow models [29, 26] to translate pretrained discrete tokens back into raw motion space. This enhances naturalness while preserving alignment with the conditioning signals.
Our method offers two primary advantages over traditional discrete decoding methods, as shown in Figure 1. First, instead of directly decoding discrete tokens into motion space, we use them as conditioning signals to guide motion generation within the continuous motion space. This reduces fine-grained noise and results in smoother, more natural motion. Second, rather than relying on a one-step decoding process, we employ an iterative refinement approach using a rectified flow model [5, 67], which progressively improves reconstruction quality. This enables the generation of dynamic and complex movements that conventional methods struggle to capture. Moreover, DisCoRD is framework-agnostic, making it adaptable to any discrete generation method (e.g., autoregressive [64] or bidirectional [12]), regardless of the conditioning signal type (e.g., text, music, or speech), thereby improving performance.
@@
Although our method improves motion naturalness, evaluating this quality remains challenging. Traditional metrics such as MPJPE do not correlate well with human perception [9, 18], and FID fails to capture subtle, frame-wise noise, as illustrated in Figure 3. These limitations reduce the reliability of existing metrics in accurately assessing reconstructed motion naturalness. To address this, we introduce a novel sample-wise metric, the symmetric Jerk Percentage Error (sJPE), which evaluates reconstructed motion by simultaneously detecting under-reconstruction and fine-grained artifacts. Our experiments demonstrate the effectiveness of DisCoRD in enhancing sample-wise naturalness without suffering from under-reconstruction. Extensive evaluations across text-to-motion, co-speech gesture, and music-to-dance generation demonstrate that our method achieves state-of-the-art naturalness, consistently outperforming existing approaches. Our contributions are as follows:
@@
    We introduce DisCoRD, a novel method for decoding discrete tokens in continuous motion space, improving the naturalness of generated motions while preserving faithfulness across various models and tasks.
@@
    We propose a novel evaluation scheme, the symmetric Jerk Percentage Error (sJPE), designed to evaluate both under-reconstruction and frame-wise noise, which are often overlooked but critical for motion generation.
@@ 2 Related Work
Human Motion Generation from Diverse Signals.
Generating natural and controllable 3D human motion remains a long-standing task. Early approaches prioritized motion naturalness [22, 2, 13], while recent advances in deep learning expand capabilities to generate motion conditioned on diverse signals. Recent advances of text-to-motion datasets [11, 42] have driven progress in text-conditioned motion synthesis, while music-motion datasets [25, 48] have enabled music-driven dance generation. Speech-motion datasets [62, 27, 28] extends motion control by synthesizing gestures from speech. As most recent methods rely on discrete representations [12, 10, 30], our work enhances motion naturalness for all discrete methods, regardless of the task, by addressing the discrete token decoding problem.
Continuous Human Motion Generation.
Early approaches to signal-to-motion generation employ regression-based mapping of control signals to motion within a continuous representation space. These works leverage Variational Autoencoders (VAE) [21, 11, 39], GANs [14], or CLIP features [43, 50] to generate natural motion. More recently, with the success of diffusion models [15, 47], motion generation models have achieved unprecedented generation quality [52, 65, 53]. Follow-up works explore continuous latent spaces for efficient motion generation [60, 8], incorporate physical constraints to improve realism [63], and integrate retrieval mechanisms to enhance generalization [66]. Their ability to generate smooth, natural motion makes them well-suited for not only motion synthesis but also for a generative prior [45, 37]. However, due to the lack of scalability in current motion datasets, the high complexity of continuous representations often makes it difficult to establish reliable cross-modal mappings between control signals and generated motions, leading to suboptimal performance.
Discrete Human Motion Generation.
Recently, to simplify the complex mapping between control signals and motions, some methods have reformulated the generation task as a discrete token classification problem, achieving notable performance in motion generation [64, 62, 46]. These approaches often employ VQ-VAEs [54] and its variants [24] to create motion tokens, which are then used to generate motion sequences via autoregressive [64, 40, 17, 62, 35, 59], or masked [12, 19, 28, 30] token prediction. More recently, discrete diffusion models have been introduced to directly denoise these discrete tokens [7, 31]. While these methods effectively bypass complex signal-to-motion mapping challenges, their inherent characteristics—such as quantization errors and discreteness result in unnatural artifacts, including under-reconstruction and frame-wise noise.
@@ 3 Method
In this section, we introduce DisCoRD, a novel method for decoding pretrained discrete representations in the raw motion domain using rectified flow. This approach enables motion generation that is both smooth and dynamic: (1) decoding in the raw motion domain preserves natural motion smoothness, and (2) utilizing rectified flow enhances expressiveness, capturing fast-paced movements. We begin by introducing rectified flow models and motion tokenization, followed by an explanation of our condition projection, conditional rectified flow decoder, and training details.
@@ 3.1 Preliminaries
Diffusion models [15, 47] have demonstrated remarkable performance due to their iterative denoising formulation, which enhances their ability to capture complex data variations and generate high-dimensional samples. However, they typically require a large number of denoising steps to produce high-quality outputs. In contrast, rectified flow [29] provides a more direct approach by framing sample generation as a transport problem, addressed through a flow matching algorithm [29, 1, 26]. Flow matching algorithm aims to construct a transport map denoted as $T:\mathbb{R}^{d}\rightarrow\mathbb{R}^{d}$, that effectively transfers observations from the source distribution $x_{0}\sim\pi_{0}$ on $\mathbb{R}^{d}$ to the target distribution $x_{1}\sim\pi_{1}$ on $\mathbb{R}^{d}$. This transport process is formalized as the following ordinary differential equation (ODE):
@@
Here, $v$ represents the vector field, and $x_{t}$ denotes the trajectory parameterized over $t\in[0,1]$. Rectified flow follows the formulation of the forward process in diffusion models [20], but its specific parameterization enables a more direct mapping between distributions and improves efficiency. Specifically, its forward process can be expressed as:
$x_{t}=tx_{1}+(1-t)x_{0},$ (2)
where $v$ is defined as $x_{1}-x_{0}$. Then, the model is trained to learn a causal approximation of $v$, denoted as $v_{\theta}$, by solving the following least squares regression problem:
@@
Once trained, samples from the target distribution $\pi_{1}$ can be generated by solving Equation (1) using an ODE solver, where the initial conditions are drawn from the source distribution $\pi_{0}$. Unlike conventional diffusion models, which require many denoising steps, rectified flow follows a nearly straight trajectory, enabling more efficient transport with much fewer denoising steps.
@@ 3.2 DisCoRD
Directly decoding discrete tokens using traditional feed-forward decoders $\mathcal{D}$ suffer from limited expressiveness and propagate token discreteness into decoded motions, resulting in under-reconstructed and noisy outputs (Figure 4). To address these issues, we propose decoding pretrained tokens in continuous space by replacing $\mathcal{D}$ with an expressive rectified flow model. Specifically, we first extract frame-wise conditioning features from discrete tokens through a Condition Projection module, and then use these features as frame-wise conditions for a Rectified Flow Decoder that synthesizes human motion from Gaussian noise. The overall pipeline of DisCoRD is depicted in Figure 2.
Condition Projection.
To enable our decoder to generate expressive motion, we first extract frame-wise conditioning features from discrete tokens $\mathbf{Z}=[\mathbf{z}_{1},\dots,\mathbf{z}_{T/q}]$. Since each token $\mathbf{z}_{t}$ encodes information spanning $q$ consecutive motion frames, we must extract $q$ distinct, frame-specific features from each token. A naïve upsampling and linear projection would result in same $q$ features from each token, and upconvolution layers would disregard the temporal correspondence between each tokens and frames. To mitigate these issues, we first repeat each token $\mathbb{z}_{t}\in\mathbb{R}^{1\times d_{code}}$ $q$ times to restore the original temporal resolution, resulting in $\mathbb{z}^{repeat}_{t}\in\mathbb{R}^{q\times d_{code}}$. Then we stack into a tensor $\mathbf{z}_{t}^{stacked}\in\mathbb{R}^{1\times(q\times d_{code})}$ and project to $\mathbb{z}^{project}_{t}\in\mathbb{R}^{1\times(q\times d_{feat})}$. Finally, we unstack the projected tensor to $\mathbb{z}^{final}_{t}\in\mathbb{R}^{q\times d_{feat}}$ where each vector $\mathbf{c}_{i}\in\mathbb{R}^{d_{feat}}(\text{for}\ i\in[1,...,q])$ are frame-wise conditioning features. This process is applied to every token in $\mathbf{Z}$, resulting in $\mathbf{C}=[\mathbf{c}_{1},\dots,\mathbf{c}_{T}]$. This approach maintains the correspondence between tokens and motion frames by explicitly extracting $q$ features from each token, ensuring that the resulting frame-wise conditioning features are well-suited for the motion decoding. Moreover, we found that our projection method enhances motion generation on unseen token sequences on stage 2.
@@
Our goal is to decode discrete tokens into natural motion while operating within a continuous space. Therefore, we do not directly map discrete tokens back into motion, but treat discrete tokens as a signal to guide motion decoding in the raw motion space. Given the frame-wise conditioning features $\mathbf{C}$ extracted from discrete tokens by the condition projection module, we train a conditional rectified flow model to reconstruct the original motion. Specifically, given a motion data distribution $P_{\mathbf{X}}$, we define the transport process from Gaussian noise $\mathbf{X_{0}}\sim\mathcal{N}(0,I)$ to motion $\mathbf{X_{1}}\sim P_{\mathbf{X}}$, guided by frame-wise conditioning features $\mathbf{C}$, formulated as:
@@
We concatenate frame-wise features $\mathbf{C}$ along the channel dimension, similar to image generation methods [67, 16], allowing each motion frame to be conditioned independently. This formulation ensures that decoding remains in the continuous space, enabling more expressive decoding. During inference, features $\mathbf{\hat{C}}$ are first extracted from tokens generated by a pretrained token generation model. This extracted features are then iteratively decoded into $\mathbf{\hat{X}_{1}}$ by solving a conditional ODE using the Euler method, progressively improving generation quality.
@@
In variants of diffusion models, training is performed over the entire data instance with fixed sequence lengths, such as 196 frames in HumanML3D [11], a standard in motion generation [51, 60]. While this approach improves reconstruction quality in stage 1, our results indicate that these improvements do not translate effectively to generation quality in stage 2. To address this limitation and improve generalization to unseen motion sequences during inference, we train our rectified flow model on sliding windows of motion frames rather than max length spans. Additionally, although conventional U-Net diffusion models often incorporate attention mechanisms to enhance performance [44], we found this strategy to be suboptimal in our context, resulting in a performance degradation.